# GLIP-OOD: Zero-Shot Graph OOD Detection with Foundation Model

### v1

@@ 3.3 Zero-shot Graph OOD Detection with ID Labels
Although the previous method is effective, it has a fundamental limitation: it requires the label names of OOD classes. Therefore, we investigate how to perform zero-shot OOD detection using only the ID label names. A simple way to achieve this is to compute the similarity scores between the query node’s embedding $g_{i}$ and only the ID label embeddings, and then normalize these ID similarity scores. These scores can serve as the output logits of an ID classifier, similar to those used in supervised graph learning methods. Therefore, we can apply any post-hoc OOD detector—such as MSP [9] or the Energy score [14]—to these ID similarity scores to perform OOD detection.
Another simple method is to use a set of negative prompts to substitute the OOD label sentences in order to capture the negation semantics with nodes [24]. Specifically, instead of comparing a node’s embedding with known OOD labels, we measure its similarity to a set of templated negative prompts such as “This node does not belong to [ID class]” or “This is not about [ID label].” These prompts are designed to express semantic exclusion and allow the model to implicitly learn what it means for a node to fall outside the ID classes. By contrasting a node’s affinity to ID label prompts and these negative prompts, we estimate its likelihood of being OOD without requiring access to the actual OOD label space. In this case, these negative prompts form the set of pseudo-OOD label texts $\mathcal{T}_{\text{OOD}}=\{\bar{T}_{1},\bar{T}_{2},\dots,\bar{T}_{K_{\text{OOD}}}\}$. Given $\mathcal{T}_{\text{ID}}$ and $\mathcal{T}_{\text{OOD}}$, we can use Eqn. 4 to compute OOD scores for any input node. However, we find that these two simple methods do not perform well in practice. Therefore, we need to design more effective approaches to enhance the GFM’s OOD awareness.
@@ 3.4 Zero-shot Graph OOD Detection with Pseudo-OOD Labels
OOD label names generation with LLM. To enhance the model’s OOD awareness, we aim to obtain negative labels that are semantically distant from the given ID labels but close to the actual, unknown OOD labels. Since all existing node-level graph OOD detection methods operate in a transductive setting, we can access the full set of nodes during training, but not their labels. Therefore, as shown in Fig. 2, we propose using LLMs to generate pseudo-OOD label names based on the initially unlabeled graph by leveraging the textual information of the nodes. We first randomly sample a small set of nodes and prompt the LLM to perform zero-shot open-world annotation. Specifically, we provide the names of the ID classes to the LLM and ask whether each unlabeled query node belongs to any of these classes, based on its textual information. The LLM is prompted to determine:
@@
If the LLM predicts that a node $v_{i}$ does not belong to any ID category, we treat it as a pseudo-OOD node. We then prompt the LLM to summarize the node’s textual content and generate a suitable OOD category name for it.
@@
OOD detection in augmented label space. Once a set of pseudo-OOD label names is generated by the LLM, we construct an augmented label space that includes both the original ID classes and the newly synthesized OOD labels. The generated OOD categories form the set $\mathcal{Y}_{\text{OOD}}$, and the augmented label space is defined as:
$\mathcal{Y}_{\text{aug}}=\mathcal{Y}_{\text{ID}}\cup\mathcal{Y}_{\text{OOD}}.$ (8)
We encode each label in $\mathcal{Y}_{\text{aug}}$ into a text embedding using the pretrained text encoder, as shown in Eqn. 2. For each node, we compute the similarity between its embedding and the embeddings of all labels in the augmented label space. A node is classified as OOD if it has low affinity to all ID label embeddings but exhibits high similarity to one or more pseudo-OOD labels. This approach allows the model to leverage both the discriminative power of ID labels and the semantic coverage of LLM-generated OOD labels, enabling more accurate and fine-grained OOD detection without requiring access to any real OOD label names.
@@ 4.1 Experimental Setup
Datasets We utilize the following TAG datasets, which are commonly used for node classification: Cora [16], Citeseer [7], Ele-Computers [30] and Wiki-CS [17]. The reason for using these datasets is that we perform zero-shot OOD detection based on GraphCLIP [33], and these datasets were not used during the pretraining of this GFM. Therefore, we can effectively evaluate its zero-shot OOD detection performance on unseen data. For each dataset, we split all classes into ID and OOD sets and the ID classes for the four datasets are shown in Appendix B. Additionally, the number of ID classes is set to a minimum of three to perform ID classification task.
Baselines We evaluate GLIP-OOD against two categories of baselines, including: (1) Supervised graph OOD detection methods, including MSP [9], Entropy, Energy [14], GNNSafe [27], and GRASP [15]. (2) Language model methods for zero-shot OOD detection, including BERT [4], SBERT [21], DeBERTa [8], and E5 [25].
@@
Settings For all methods that require labeled nodes to train the ID classifier, we follow the setting in [22] and use $20\times K$ labeled ID nodes for training on each dataset with $K$ ID classes. Additionally, we randomly select $10\times K$ labeled ID nodes and an equal number of OOD nodes for validation. The test set consists of 500 randomly selected ID nodes and 500 OOD nodes. For language model-based methods, we use them to encode the text associated with nodes and the information of ID labels, respectively. We then compute similarity scores between each node embedding and the embeddings of ID label sentences. Based on the normalized similarity scores, we calculate the MSP score, which is used as the OOD score. For GLIP-OOD, we use GPT-4o-mini to identify pseudo-OOD nodes and generate corresponding pseudo-OOD label names. The prompts used in this process are provided in Appendix D. Our method has two variants: (1) using all real labels, and (2) using both ID labels and pseudo-OOD labels to perform zero-shot OOD detection.
Implementation details For all supervised methods, we employ a two-layer GCN as the ID classifier and apply various post-hoc OOD detectors to it. Each GCN is configured with a hidden dimension of 32, a learning rate of 0.01, a dropout rate of 0.5 and a weight decay of 5e-4. For all experiments, results are averaged over 5 different random seeds. All experiments are conducted on hardware equipped with an NVIDIA RTX 4080 Ti.
@@
Evaluation Metrics For the ID classification task, we use classification accuracy (ACC) as the evaluation metric. For the OOD detection task, we employ three commonly used metrics from the OOD detection literature [27]: the area under the ROC curve (AUROC), the precision-recall curve (AUPR), and the false positive rate when the true positive rate reaches 95% (FPR@95). In all experiments, OOD nodes are considered positive cases.
@@ 4.2 Main Results
The OOD detection results are shown in Table 1. From the results, we conclude the following: (1) Exposing all class labels to GLIP-OOD enables it to achieve the best performance among all methods, even surpassing fully supervised approaches that rely on a large number of ID labels to train the ID classifier. This demonstrates that the GFM is inherently strong enough to perform zero-shot OOD detection and has significant potential to become highly OOD-aware, provided the OOD classes are clearly defined. (2) For language model baselines performing zero-shot OOD detection, relying solely on ID class labels is inadequate for distinguishing ID and OOD nodes in most cases. (3) Using our proposed approach to generate pseudo-OOD labels significantly improves the OOD awareness of the GFM, compared to using only ID class labels to compute similarity scores and applying post-hoc OOD detector on top of the language model’s output logits. (4) When using ID labels and pseudo-OOD labels, GLIP-OOD achieves performance comparable to that of supervised methods. However, GLIP-OOD requires neither labeled nodes for training nor any knowledge of OOD labels. The only information it needs is the names of the ID classes, after which it can directly perform zero-shot OOD detection for any input node. (5) While GLIP-OOD is primarily designed for OOD detection, it also performs competitively in ID classification. Although it lags slightly behind fully supervised GNN-based classifiers in terms of accuracy (as expected given the absence of node-level supervision), its performance remains strong enough to support reliable OOD detection.
Table 1: Performance comparison (best OOD detection results highlighted in bold) of different models on ID classification and OOD detection tasks.
@@
Note: GLIP-OOD-L denotes the use of labels generated by the LLM, GLIP-OOD-R denotes the use of real labels.
@@ 4.3 Visualization of Label Space
Figure 3: Visualization of the semantic label space for the Ele-Computers and Wiki-CS dataset. Each point represents a label name embedded in the GLIP-OOD’s semantic space.
@@ 4.4 More OOD Scores
We use additional OOD scoring functions for zero-shot OOD detection with all labels. Following Eqn. 5, the other three OOD scores are defined as follows:
@@
From the results in Table 2, we observe that all four post-hoc OOD scores generally yield similar OOD detection performance, and GLIP-OOD remains robust across different OOD scoring functions.
@@ 4.5 Zero-shot OOD Detection with Weak OOD Information
As shown in Table 3, GLIP-OOD consistently outperforms all language models across all datasets. This highlights the effectiveness of the GFM; even without training on node labels, the GFM can effectively leverage label names and graph structure to learn semantically meaningful representations, enabling robust OOD detection in a fully zero-shot setting.
@@ 4.6 More OOD and ID Class Divisions
GLIP-OOD 0.9273 0.9185 0.3484 0.9253 0.9182 0.7016 0.8919 0.8975 0.4952 0.9328 0.9299 0.3008
To further evaluate the robustness and generalizability of GLIP-OOD, we experiment with an alternative set of OOD and ID class divisions across four datasets. The new ID classes are listed in Appendix B. From the results in Table 4, we observe that GLIP-OOD consistently achieves the best performance across all datasets and metrics. This demonstrates that the GFM possesses strong inherent capability for OOD detection and can adapt to different open-set configurations, thereby enhancing both the reliability and robustness of zero-shot graph OOD detection.
@@ 5.2 Zero-shot OOD Detection
Zero-shot OOD detection has been widely studied in the image domain [24, 5, 18, 1] and text domain [32, 10]. These methods typically provide large-scale pretrained models with only the label names of ID classes and use similarity scores to distinguish ID from OOD inputs. For example, EOE [1] designs LLM prompts guided by visual similarity to generate potential outlier class labels tailored for zero-shot OOD detection. ZOC [5] extends the CLIP model to dynamically generate candidate unseen labels for each test image and introduces a novel confidence score based on the similarity between the test image and both seen and generated unseen labels in the feature space. CLIPN [24] enhances CLIP with the ability to distinguish between ID and OOD samples by leveraging positive-semantic and negation-semantic prompts. [10] proposes an ID-label-free OOD detection method that leverages unsupervised clustering and contrastive learning to learn effective data representations for OOD detection. Despite these advances, zero-shot OOD detection in graph domains remains largely unexplored. This is due to the unique challenges posed by relational inductive biases and the lack of large-scale, general-purpose foundation models specifically trained for graphs. Our work bridges this gap by enabling zero-shot node-level OOD detection in graphs through the use of GFMs and language models for pseudo-OOD label generation—without requiring any labeled nodes.
@@ 6 Conclusion
In this work, we introduce GLIP-OOD, a novel framework for zero-shot graph OOD detection that leverages the capabilities of graph foundation model. Unlike existing graph OOD detection methods that rely on large numbers of labeled ID nodes, GLIP-OOD operates in a fully zero-shot setting, requiring only the names of ID classes and the graph structure. By prompting LLMs to generate pseudo-OOD labels, our method enables GFMs to learn meaningful semantic distinctions between ID and OOD nodes without any supervised training. Through extensive evaluation on four benchmark datasets, we demonstrate that when provided with all label names, GLIP-OOD outperforms all supervised methods. When given only ID label names, it achieves performance comparable to supervised approaches that rely on a large number of labeled ID nodes for training. Future research could explore more effective strategies for generating negative labels to further enhance the inherent OOD-awareness of increasingly powerful graph foundation models.
Broader Impact By enhancing the OOD-awareness of graph-based systems, our method advances the development of safer and more trustworthy AI models for open-world environments. It also paves the way for future research at the intersection of foundation models, LLMs, and graph OOD detection, offering a promising step toward scalable, reliable, and label-efficient open-world graph learning.
Limitations While our proposed method eliminates the need for both labeled nodes and a training process for graph OOD detection, it is applicable only to TAGs. Additionally, reliance on LLMs for label generation introduces concerns regarding the biases present in the pretraining of those models.

### v2

@@ 3.3 Zero-shot Graph OOD Detection with ID Labels
Although the previous method is effective, it has a fundamental limitation: it requires the label names of OOD classes. Therefore, we investigate how to perform zero-shot OOD detection using only the ID label names. A simple way to achieve this is to compute the similarity scores between the query node’s embedding $g_{i}$ and only the ID label embeddings, and then normalize these ID similarity scores. These scores can serve as the output logits of an ID classifier, similar to those used in supervised graph learning methods. Therefore, we can apply any post-hoc OOD detector—such as MSP [9] or the Energy score [14]—to these ID similarity scores to perform OOD detection. However, in this setting, the model has no explicit notion of what constitutes an OOD class; it can only identify OOD nodes based on their weak affinity to ID labels.
To mitigate this limitation, a straightforward approach is to use a set of negative prompts that express semantic exclusion—such as “This node does not belong to [ID class]” or “This is not about [ID label]” [26]—to serve as pseudo-OOD label sentences. These prompts aim to capture negation semantics and enable the model to infer OOD-ness by contrasting a node’s similarity to ID labels versus these negative prompts. Formally, the set of negative prompts serves as the pseudo-OOD label set $\mathcal{T}_{\text{OOD}}=\{\bar{T}_{1},\bar{T}_{2},\dots,\bar{T}_{K_{\text{OOD}}}\}$, which, together with $\mathcal{T}_{\text{ID}}$, is used in Eqn. 4 to compute OOD scores. However, we observe that this approach underperforms in practice. Negative prompts tend to introduce ambiguous or weak semantic signals, making it difficult for the model to reliably distinguish OOD nodes—especially when OOD cases are semantically close to ID classes. This motivates the need for more effective strategies to enhance the GFM’s OOD awareness.
@@ 3.4 Zero-shot Graph OOD Detection with Pseudo-OOD Labels
OOD label names generation with LLM. To enhance the model’s OOD awareness, we aim to obtain negative labels that are semantically distant from the given ID labels but close to the actual, unknown OOD labels. Unlike image and text OOD detection, node-level graph OOD detection typically operates in a transductive setting, where the full set of nodes is accessible during training, but only the labels for ID classes are provided. Since all existing methods follow this setting, we propose leveraging LLMs to generate pseudo-OOD label names based on the initially unlabeled graph. As illustrated in Fig. 2, our approach utilizes the textual information associated with nodes to prompt the LLM to infer semantically meaningful pseudo-OOD labels without requiring access to real OOD data. We begin by randomly sampling a small set of unlabeled nodes and prompting the LLM to perform zero-shot open-world annotation. Specifically, we provide the names of the ID classes to the LLM and ask whether each unlabeled query node belongs to any of these ID classes, based on its textual information. The LLM is prompted to determine:
@@
If the LLM predicts that a node $v_{i}$ does not belong to any ID class, we treat it as a pseudo-OOD node. We then prompt the LLM to summarize the node’s textual content and generate a suitable OOD category name for it:
@@
OOD detection in augmented label space. Once a set of pseudo-OOD label names is generated by the LLM, we construct an augmented label space that includes both the original ID labels and the newly synthesized OOD labels. The generated OOD labels form the set $\mathcal{Y}_{\text{OOD}}$, and the augmented label space is defined as $\mathcal{Y}_{\text{aug}}=\mathcal{Y}_{\text{ID}}\cup\mathcal{Y}_{\text{OOD}}$. From $\mathcal{Y}_{\text{aug}}$, we obtain the corresponding label sentences $\mathcal{T}_{\text{aug}}$, each of which is encoded into a text embedding using the pretrained text encoder, as shown in Eqn. 2. During inference, for each node, we compute its similarity to all label embeddings in the augmented space using Eqn. 4. A node is classified as OOD if it has low affinity to all ID label embeddings but exhibits high similarity to one or more pseudo-OOD label embeddings. This enables the model to leverage both the discriminative power of ID labels and the semantic richness of LLM-generated OOD labels, facilitating accurate and fine-grained OOD detection without requiring access to real OOD label names.
@@ 4.1 Experimental Setup
Datasets. We utilize the following TAG datasets, which are commonly used for node classification: Cora [17], Citeseer [7], Ele-Computers [32], and Wiki-CS [18] (see Appendix A for details). We use these datasets because we perform zero-shot OOD detection based on GraphCLIP [35], and these datasets were not used during the pretraining of this GFM. Therefore, we can effectively evaluate its zero-shot OOD detection performance on unseen data. For each dataset, we split all classes into ID and OOD sets and the ID classes for the four datasets are shown in Appendix B. Additionally, the number of ID classes is set to a minimum of three to perform the ID classification task.
Baselines. We evaluate GLIP-OOD against two categories of baselines. (1) GNN-based OOD detection with ID supervision, which trains a GNN classifier using labeled ID nodes and then applies post-hoc scoring functions to the output logits for OOD detection. This includes MSP [9], Entropy, Energy [14], GNNSafe [29], and GRASP [16]. (2) Language model-based zero-shot OOD detection, which computes similarity between textual node features and ID class names without using any labeled nodes. This includes BERT [4], SBERT [23], DeBERTa [8], and E5 [27]. Here, we use only discriminative language models for OOD detection, as generative LLMs (e.g., GPT) cannot produce soft OOD scores (e.g., those defined in Eqn. 4) from their output logits; instead, they yield only hard binary predictions (e.g., yes or no).
@@
Settings. For all methods that require labeled nodes to train the ID classifier, we follow the setting in [24] and use $20\times K$ labeled ID nodes for training on each dataset with $K$ ID classes. Additionally, we randomly select $10\times K$ labeled ID nodes and an equal number of OOD nodes for validation. The test set consists of 500 randomly selected ID nodes and 500 OOD nodes. For language model-based methods, we use them to encode the text associated with nodes and the information of ID labels, respectively. We then compute similarity scores between each node embedding and the embeddings of ID label sentences. Based on the normalized similarity scores, we calculate the MSP, which is used as the OOD score (results using other OOD scores are provided in Appendix C). Our method has two variants: (1) using all real labels, and (2) using both ID labels and pseudo-OOD labels to perform zero-shot OOD detection.
Implementation details. For all supervised methods, we employ a two-layer GCN as the ID classifier and apply various post-hoc OOD detectors to it. Each GCN is configured with a hidden dimension of 32, a learning rate of 0.01, a dropout rate of 0.5 and a weight decay of 5e-4. For GLIP-OOD, we use GPT-4o-mini to identify pseudo-OOD nodes and generate pseudo-OOD label names. In our experiments, we randomly sample 80 nodes from each of the Cora and Citeseer datasets, and 120 nodes from each of the Ele-Computers and Wiki-CS datasets. The LLM is prompted to classify each node into one of the predefined ID classes or, if it determines that the node does not belong to any ID class, to generate a new pseudo-OOD label name. During zero-shot inference, we randomly select a subset of the generated labels to construct the pseudo-OOD label set, and perform OOD detection using both the pseudo-OOD and original ID labels. Further implementation details and the prompts used in this process are provided in Appendix D. All results are averaged over 5 random seeds, and all experiments are conducted on a machine equipped with an NVIDIA GeForce RTX 4080 SUPER.
@@
Evaluation Metrics. For the ID classification task, we use classification accuracy (ACC) as the evaluation metric. For the OOD detection task, we employ three commonly used metrics from the OOD detection literature [29]: the area under the ROC curve (AUROC), the precision-recall curve (AUPR), and the false positive rate when the true positive rate reaches 95% (FPR@95). In all experiments, OOD nodes are considered positive cases.
@@ 4.2 Main Results
The OOD detection results on AUROC, AUPR and FRP@95 are shown in Table 1. From the results, we conclude the following: (1) Access to all class labels yields state-of-the-art performance. Exposing all class labels to GLIP-OOD enables it to achieve the best performance among all methods, even surpassing supervised approaches that rely on a large number of ID labeled nodes to train the ID classifier. This demonstrates the strong inherent capability of the GFM to perform zero-shot OOD detection. When provided with well-defined class semantics, the GFM can distinguish ID from OOD nodes without fine-tuning, highlighting its high potential for OOD awareness. (2) Using only ID labels limits zero-shot OOD detection. For language model baselines performing zero-shot OOD detection, relying solely on ID class labels is generally inadequate for distinguishing between ID and OOD nodes. This highlights the need for either additional supervision or synthetic augmentation to more effectively capture the semantic boundary between ID and OOD classes. (3) Pseudo-OOD label generation enhances OOD awareness. The introduction of pseudo-OOD labels generated by our proposed approach significantly enhances the OOD awareness of the GFM. By constructing a contrastive label space that incorporates synthetic OOD semantics, our method improves the model’s ability to calibrate OOD scores and more effectively distinguish ID from OOD nodes. This approach outperforms post-hoc OOD detection methods applied to the output logits of language models. (4) GLIP-OOD is competitive with supervised baselines—without supervision. When using ID labels and pseudo-OOD labels, GLIP-OOD achieves performance comparable to that of supervised methods. However, GLIP-OOD requires neither labeled nodes for training nor any knowledge of OOD labels. The only information it needs is the names of the ID classes, after which it can directly perform zero-shot OOD detection for any input node. (5) Moderate Zero-shot ID Classification Performance. While GLIP-OOD is primarily designed for OOD detection, it also performs competitively in ID classification. Although it lags slightly behind fully supervised GNN-based classifiers in terms of accuracy (ACC shown in Table 1)—as expected due to the absence of node-level supervision—its performance remains strong enough to support reliable OOD detection.
Table 1: Performance comparison (best OOD detection results highlighted in bold) of different models on ID classification and OOD detection tasks. GLIP-OOD-L denotes the use of pseudo-labels generated by the LLM, while GLIP-OOD-R denotes the use of real labels.
@@ 4.3 Visualization of Label Space
Figure 3: Visualization of the semantic label space. Each point corresponds to a class label name embedded by GLIP-OOD (green for ID labels, purple for real OOD labels, and blue for pseudo-OOD labels by the LLM). The pseudo-OOD labels lie outside the ID cluster but remain semantically close to real OOD labels, highlighting the effectiveness as proxies for enhancing OOD awareness.
@@ 4.4 More OOD Scores
We use additional OOD scoring functions for zero-shot OOD detection with all labels. Following Eqn. 5, the other three OOD scores are defined in Eqn. 4.4. From the results in Table 2, we observe that all four post-hoc OOD scores generally yield similar OOD detection performance, and GLIP-OOD remains robust across different OOD scoring functions.
@@ 4.5 Zero-shot OOD Detection with Weak OOD Information
As shown in Table 3, GLIP-OOD consistently outperforms all language models across all datasets. This highlights the effectiveness of the GFM; even without training on node labels, it can leverage label names and graph structure to learn semantically meaningful representations, enabling robust OOD detection in a fully zero-shot setting.
@@ 4.6 More OOD and ID Class Divisions
GLIP-OOD-R 0.9273 0.9185 0.3484 0.9253 0.9182 0.7016 0.8919 0.8975 0.4952 0.9328 0.9299 0.3008
To further evaluate the robustness and generalizability of GLIP-OOD, we experiment with an alternative set of OOD and ID class divisions across four datasets. The new ID classes are listed in Appendix B. From the results in Table 4, we observe that GLIP-OOD consistently achieves the best performance across all datasets and metrics. This demonstrates that the GFM possesses strong inherent capability for OOD detection and can adapt to different open-set configurations, thereby enhancing the reliability and robustness of zero-shot graph OOD detection.
@@ 5.2 Zero-shot OOD Detection
Zero-shot OOD detection has been widely studied in the image domain [26, 5, 19, 1, 15] and text domain [34, 10]. These methods typically provide large-scale pretrained models with only the label names of ID classes and use similarity scores to distinguish ID from OOD inputs. For example, EOE [1] designs LLM prompts guided by visual similarity to generate potential outlier class labels tailored for zero-shot OOD detection. ZOC [5] extends the CLIP model to dynamically generate candidate unseen labels for each test image and introduces a novel confidence score based on the similarity between the test image and both seen and generated unseen labels in the feature space. CLIPN [26] enhances CLIP with the ability to distinguish between ID and OOD samples by leveraging positive-semantic and negation-semantic prompts. [10] proposes an ID-label-free OOD detection method that leverages unsupervised clustering and contrastive learning to learn effective data representations for OOD detection. Despite these advances, zero-shot OOD detection in graph domains remains largely unexplored. This is due to the unique challenges posed by relational inductive biases and the lack of large-scale, general-purpose foundation models specifically trained for graphs. Our work bridges this gap by enabling zero-shot node-level OOD detection in graphs through the use of GFMs and language models for pseudo-OOD label generation—without requiring any labeled nodes.
@@ 6 Conclusion
In this work, we introduce GLIP-OOD, a novel framework for zero-shot graph OOD detection that leverages the capabilities of graph foundation model. Unlike existing graph OOD detection methods that rely on large numbers of labeled ID nodes, GLIP-OOD operates in a fully zero-shot setting, requiring only the names of ID classes and the graph structure. By prompting LLMs to generate pseudo-OOD labels, our method enables GFMs to learn meaningful semantic distinctions between ID and OOD nodes without any supervised training. Through an extensive evaluation on four benchmark datasets, we demonstrate that when provided with all label names, GLIP-OOD outperforms all supervised methods. When given only ID label names, it achieves performance comparable to supervised approaches that rely on a large number of labeled ID nodes for training. Future research could explore more effective strategies for generating negative labels to further enhance the inherent OOD-awareness of increasingly powerful graph foundation models.
Broader Impact. By enhancing the OOD-awareness of graph-based systems, our method advances the development of safer and more trustworthy AI models for open-world environments. It also paves the way for future research at the intersection of foundation models, LLMs, and graph OOD detection, offering a promising step toward scalable, reliable, and label-efficient open-world graph learning.
Limitations. While our proposed method eliminates the need for both labeled nodes and a training process for graph OOD detection, it is applicable only to TAGs. Additionally, reliance on LLMs for label generation introduces concerns regarding the biases present in the pretraining of those models.
Appendix: GLIP-OOD: Zero-Shot Graph OOD Detection with Graph Foundation Model